OpenStudy (anonymous):

Can someone check my answers? Use the Change of Base Formula to evaluate log4 27. Then convert log4 27 to a logarithm in base 2. Round to the nearest thousandth. I got 2.377; log_2 5.196

OpenStudy (anonymous):

1 is good 2 is not

OpenStudy (misty1212):

\[\log_b(A)=\frac{\ln(A)}{\ln(b)}\]

OpenStudy (anonymous):

2.377 is good but 5.196 is not. i think they want to know what log[2] 27 is

OpenStudy (anonymous):

the 2nd one is log_2 13.5, right?

OpenStudy (anonymous):

no, you want to find log_2 27. remember that 4 = 2^2

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

I have another question

OpenStudy (anonymous):

did you want to finish this one first?

OpenStudy (anonymous):

Choose the property used to rewrite the expression log 15sqrt(125x^3) = 1/5 log5 x I put Power Property

OpenStudy (anonymous):

I got it for the first one

OpenStudy (anonymous):

I had to round it at the end